3. Use De Moivre’s Theorem to determine the cube root of Z and leave your answer in polar


form with the angle in radians


(a) Z = 1+i√3


1
Expert's answer
2022-05-12T09:23:25-0400

At first, we rewrite "z=1+i\sqrt{3}" in the polar form. We get: "z=2(\frac12+i\frac{\sqrt{3}}2)=2(\cos(\frac{\pi}{3})+i\,\sin(\frac{\pi}{3}))".

We use the known formula (extension of de Moivre's formula) to compute the root "\frac13". We get:

"z^{\frac13}=2^{\frac13}(\cos(\frac{\frac{\pi}3+2\pi k}{3})+i\,\sin(\frac{\frac{\pi}3+2\pi k}{3}))," where "k=0,1,2."

We receive the following roots:

"v_0=2^{\frac13}(\cos(\frac{{\pi}}{9})+i\,\sin(\frac{{\pi}}{9}))",

"v_1=2^{\frac13}(\cos(\frac{{7\pi}}{9})+i\,\sin(\frac{{7\pi}}{9})),"

"v_2=2^{\frac13}(-\cos(\frac{{4\pi}}{9})-i\,\sin(\frac{{4\pi}}{9}))".

Answer: the cube root of "z" has the following values:

"v_0=2^{\frac13}(\cos(\frac{{\pi}}{9})+i\,\sin(\frac{{\pi}}{9}))",

"v_1=2^{\frac13}(\cos(\frac{{7\pi}}{9})+i\,\sin(\frac{{7\pi}}{9})),"

"v_2=2^{\frac13}(-\cos(\frac{{4\pi}}{9})-i\,\sin(\frac{{4\pi}}{9}))".